OpenStudy (anonymous):

2^0 +3 .... does that make it 2^1 +3 which = 5 or does it make it 1+3=4 ?

OpenStudy (jamesj):

We take 2^0 to be equal to 1.

OpenStudy (anonymous):

okay so even though the 2 is there it just equals one due to the fact that it was a 0?

OpenStudy (jamesj):

Yes. In general, for all positive numbers a, a^0 = 1. Which makes sense in the rules of exponents because \[a^0 \times a^1 = a^{0+1} = a^1 \] hence we should expect that a^0 is the multiplicative identity, i.e., the number 1.

OpenStudy (anonymous):

okay! thanks so much!